OpenStudy (anonymous):

Write the point-slope form of the equation passing through (5, -1) with a slope of 6.

OpenStudy (lυἶცἶ0210):

Have you tried using the point-slope formula? \(\large y-y_{1}=m(x-x_{1}) \) ?

OpenStudy (anonymous):

yes

OpenStudy (lυἶცἶ0210):

What do you need help with specifically? Plugging in the information?

OpenStudy (anonymous):

yea

OpenStudy (lυἶცἶ0210):

Relating to the equation: \(m=slope \) and the coordinates (5, -1) are \((x_{1}, y_{1})\)

OpenStudy (anonymous):

okay

OpenStudy (lυἶცἶ0210):

So just plug them in.. \(\Large y-(-1)=6(x-5) \)

OpenStudy (anonymous):

Oh i got it thanks so much (:

OpenStudy (lυἶცἶ0210):

You're welcome :) And welcome to OpenStudy btw \(\Huge \color{green}{\star^{\star^{\color{red}{\star}}}} \)